OpenStudy (anonymous):

BEST RESPONSE graph x = 3 sin^3 t y = 3 cos^3 t

OpenStudy (goformit100):

yes why am i tagged here ?

OpenStudy (anonymous):

are you going to answer my question?

OpenStudy (anonymous):

one can show (I can do the proof if you'd like to see it) that for x, y >= 0 what you described is equivalent to the equation \[(\frac{x}{3})^{\frac{2}{3}} + (\frac{y}{3})^{\frac{2}{3}} = 1\]This might help you to understand how to plot this. For example, consider the case x = 3. It follows that y = 0. For y = 3, it follows that x = 0. The bigger x is, the smaller y must be (direct impliciation of the equation), so we can deduce that the curve will be falling down smoothly on the way from (0,3) to (3,0).